OpenStudy (anonymous):

Simplify. x^8 ÷ x^4 · x^2 a) 1 b) x2 c) x4 d) x6

OpenStudy (anonymous):

d

OpenStudy (anonymous):

here's why :

OpenStudy (anonymous):

x^8 / x^4 = x^(8-4) * x^2 = x^(4+2) = x^6. D it is.

OpenStudy (anonymous):

ya, what TBates said, haha